OpenStudy (anonymous):

Irene is mixing starch with flour to make a pinata. For every cup of flour, she needs to add 1/4 cup of starch. How much starch does Irene need for 10 cups of flour? Express your answer in simplest form. cups

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

please

OpenStudy (anonymous):

kelliegirl33??

OpenStudy (anonymous):

are you there

OpenStudy (anonymous):

your sis asked me to come here :) proportion time... 1/(1/4) = 10/x (1 cup flour to 1/4 cup starch = 10 cups flour to x cups starch) cross multiply (1)(x) = (1/4)(10) x = 10/4 x = 2 1/2 cups starch ==> your answer

OpenStudy (anonymous):

ok i have 2 more

OpenStudy (anonymous):

are good with that

OpenStudy (anonymous):

lets see what you got

OpenStudy (anonymous):

A whale's heart can beat once every 1/9 of a minute when it is resting. How many times can the whale's heart beat in 5 minutes?

OpenStudy (anonymous):

hello??

OpenStudy (anonymous):

are you done yet

OpenStudy (anonymous):

Is this multiple choice ?

OpenStudy (anonymous):

hello??

OpenStudy (anonymous):

is this multiply choice ?

OpenStudy (anonymous):

no you have to pick a number no choices

OpenStudy (anonymous):

1/9 per minute means 5/9 in 5 minutes. 5/9 of a minute = 5/9 * 60 = 300/9 = 33.33 sec. 33.33 x 5 = 166.65 rounds to 167 beats in 5 minutes ?? You better get a second opinion on this one

OpenStudy (anonymous):

Michael swims back and forth across his pond. It is 1/4 miles across the pond. How many times will Michael need to swim across his pond to swim 8 miles? A. 32 B. 16 C. 8 D. 2

OpenStudy (anonymous):

8/(1/4) = 8 * 4/1 = 32 check.. 32(1/4) = 8 32/4 = 8 8 = 8 (correct) answer is 32 times

OpenStudy (anonymous):

got anymore ?

OpenStudy (anonymous):

i didint past 77%

OpenStudy (anonymous):

77 is passing

OpenStudy (anonymous):

no i have to get 80 or more

OpenStudy (anonymous):

that sucks....I am sorry